Find the first moment of area under the curve y=1+x+x^2,from x=0,x=2.
Answer by ewatrrr(24785)

y=1+x+x^2
find Integral:
f(x) = x^3/3 + x^2/2 + x
Evaluate f(2) - f(0)
8/3 + 2 + 2 = 12/3 + 8/3 = 20/3 = 6 2/3